### v1

# Inverse Problem Sampling in Latent Space Using Sequential Monte Carlo

Abstract
In image processing, solving inverse problems is the task of finding plausible reconstructions of an image that was corrupted by some (usually known) degradation model. Commonly, this process is done using a generative image model that can guide the reconstruction towards solutions that appear natural. The success of diffusion models over the last few years has made them a leading candidate for this task. However, the sequential nature of diffusion models makes this conditional sampling process challenging. Furthermore, since diffusion models are often defined in the latent space of an autoencoder, the encoder-decoder transformations introduce additional difficulties. Here, we suggest a novel sampling method based on sequential Monte Carlo (SMC) in the latent space of diffusion models. We use the forward process of the diffusion model to add additional auxiliary observations and then perform an SMC sampling as part of the backward process. Empirical evaluations on ImageNet and FFHQ show the benefits of our approach over competing methods on various inverse problem tasks.
Machine Learning, ICML
1 Introduction
Many important signal processing tasks can be viewed as inverse problems (Song et al., 2021b; Moliner et al., 2023; Daras et al., 2024; Chung et al., 2023b; Cardoso et al., 2023). In inverse problems, the objective is to obtain a clean signal ${\mathbf{x}}\in\mathbb{R}^{n}$ from a degraded observation ${\mathbf{y}}={\mathcal{A}}({\mathbf{x}})+\bm{\mathbf{\psi}}$, where ${\mathcal{A}}$ is usually a known irreversible mapping and $\bm{\mathbf{\psi}}$ is a Gaussian noise vector. Common applications that fit this framework include image deblurring, super-resolution, inpainting, and Gaussian denoising. The broad applicability of inverse problems makes them highly significant, as they encompass numerous real-world challenges, such as those found in digital image processing (Blackledge, 2005), wireless communication (Chen et al., 2021), seismology (Virieux & Operto, 2009), medical imaging (Song et al., 2021b; Chung et al., 2023c), and astronomy (Craig & Brown, 1986).
A major challenge in solving inverse problems is the existence of multiple plausible solutions. For example, in image inpainting, the likelihood $p({\mathbf{y}}|{\mathbf{x}})$ remains constant regardless of how the absent pixels are filled. However, the desired solution is one that not only fits the observation, but also appears natural, which corresponds to having a high probability under a natural image prior $p({\mathbf{x}})$. This insight naturally leads to the approach of sampling from the posterior distribution $p({\mathbf{x}}|{\mathbf{y}})\propto p({\mathbf{y}}|{\mathbf{x}})p({\mathbf{x}})$, combining the data likelihood and the prior to achieve realistic and data-consistent solutions.
Figure 1: LD-SMC solves inverse problems in the latent space by first augmenting the model with auxiliary observations using the forward diffusion process. Then, sampling from the posterior distribution $p_{\theta}({\mathbf{z}}_{0}|{\mathbf{y}}_{0})$ is done based on the backward diffusion process using sequential Monte Carlo. In the figure, ${\mathcal{D}}$ and ${\mathcal{A}}$ denote the decoder and the corruption operator respectively.
With the impressive recent advances in diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021a), there has been a significant interest in leveraging them as prior image models to solve inverse problems. However, integrating diffusion models into this context is not straightforward because of their sequential sampling process. Specifically, diffusion sampling involves iterative drawing from $p({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t})$, while the conditioning on the corrupted image ${\mathbf{y}}$ is defined only in the final step, namely, through $p({\mathbf{y}}|{\mathbf{x}}_{0})$. This mismatch makes direct sampling from the joint posterior $p({\mathbf{x}}_{0},...,{\mathbf{x}}_{T}|{\mathbf{y}})$ particularly challenging.
Figure 2: Comparison between LD-SMC and baseline methods on inpainting of images from the ImageNet dataset.
A simple solution proposed in (Chung et al., 2023b) is to approximate $p({\mathbf{y}}|{\mathbf{x}}_{t})$ with $p({\mathbf{y}}|\mathbb{E}[{\mathbf{x}}_{0}|{\mathbf{x}}_{t}])$ for efficient sampling. Further work in (Wu et al., 2024) applied a sequential Monte Carlo (SMC) process to correct for this approximation. Another recent approach proposed in (Dou & Song, 2024) connected ${\mathbf{x}}_{t}$ to ${\mathbf{y}}$ by introducing a sequence of latent variables ${\mathbf{y}}_{t}$ and sampling sequentially from $p({\mathbf{x}}_{t:T}|{\mathbf{y}}_{t:T})$. While this approach has shown great potential, it is limited to linear corruption models. As such, it is not applicable for non-linear mappings ${\mathcal{A}}$ or common Latent Diffusion Models (LDMs) (Rombach et al., 2022) due to the nonlinearity of the decoder. This is a serious restriction, as many of the recent powerful and efficient models are LDMs (Esser et al., 2024).
Both existing approaches have pros and cons. Using the $p({\mathbf{y}}|\mathbb{E}[{\mathbf{x}}_{0}|{\mathbf{x}}_{t}])$ approximation can be helpful in capturing the large scale semantics of the image, but it might not be well suited for capturing the small details. On the other hand, the auxiliary ${\mathbf{y}}_{1:T}$ can aid in the finer details and could be used with LDMs. Here, we propose a method that combines these two approaches and strives to achieve the best of both worlds. Specifically, we augment the model with additional latent variables ${\mathbf{y}}_{t}$, one for each time step, and then apply posterior inference over the latent diffusion variables ${\mathbf{z}}_{t}$. To obtain a tractable sampling procedure, we derive a novel posterior approximation and define a new proposal distribution for the SMC sampling process (Doucet et al., 2001; Del Moral et al., 2012). Hence, we name our method Latent Diffusion Sequential Monte Carlo, or more concisely LD-SMC. Importantly, as our approach utilizes these auxiliary variables in the sampling process, we found that it was better suited to more challenging inverse problems such as inpainting. An illustration of our approach is shown in Figure 1.
We empirically validated our approach on the ImageNet (Russakovsky et al., 2015) and FFHQ (Karras et al., 2019) datasets. LD-SMC usually outperforms or is comparable to baseline methods on image deblurring and super-resolution tasks, and can significantly improve over baseline methods on inapainting tasks, especially on ImageNet which has more diversity in it.
To conclude, in this study we make the following contributions: (1) A novel SMC procedure for solving inverse problems using latent diffusion models; (2) A novel posterior approximation and proposal distribution to perform approximate posterior sampling; (3) LD-SMC outperforms baseline methods, especially on challenging inpainting tasks.
2 Background
Inverse Problems. In inverse problems one would like to recover a sample ${\mathbf{x}}\in{\mathbb{R}}^{n}$ from a corrupted version of it ${\mathbf{y}}\in{\mathbb{R}}^{m}$. Usually, the corruption model that acted on ${\mathbf{x}}$ is known, but the operation is irreversible (Tarantola, 2005). For instance, restoring a high-quality image from a low-quality one. We denote the corruption operator by ${\mathcal{A}}(\cdot)$, and assume that ${\mathbf{y}}={\mathcal{A}}({\mathbf{x}})+\bm{\mathbf{\psi}}$, where $\bm{\mathbf{\psi}}\sim\mathcal{N}(0,\tau^{2}{\mathbf{I}})$ has a known standard deviation $\tau$. In a more concise way, $p({\mathbf{y}}|{\mathbf{x}})=\mathcal{N}({\mathcal{A}}({\mathbf{x}}),\tau^{2}{\mathbf{I}})$. Common examples of inverse problems are inpainting, colorization, and deblurring. In general, solving inverse problem tasks is considered an ill-posed problem with many possible solutions ${\mathbf{x}}$ with equally high $p({\mathbf{y}}|{\mathbf{x}})$ values. Given a prior distribution $p({\mathbf{x}})$ over natural images, one standard approach to solving the inverse problem is to sample the posterior distribution $p({\mathbf{x}}|{\mathbf{y}})\propto p({\mathbf{y}}|{\mathbf{x}})p({\mathbf{x}})$.
Diffusion Models. Owing to their high-quality generation capabilities, in recent years diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) have been leveraged as priors in inverse problems (Jalal et al., 2021; Song et al., 2021c). Here, we adopt the DDIM formulation (Song et al., 2021a) for the prior model, although our approach can work with other diffusion model formulations as well. Furthermore, since it is costly to apply the diffusion process in the pixel space, a common approach is to apply the diffusion model in the latent space given by an auto-encoder (Rombach et al., 2022). Applying diffusion models in the latent space allows us to sample high-quality images while reducing the computational resources needed by the model. Hence, designing models that effectively solve inverse problems using latent diffusion models is of great importance.
Denote by ${\mathbf{z}}_{1:T}$ the random variables in the latent space. Let $\alpha_{1:T}$, $\beta_{1:T}$ be the variance schedule of the diffusion process with $\beta_{t}\coloneqq 1-\alpha_{t}$. Also, denote by $\bar{\alpha}_{t}\coloneqq\prod_{j=1}^{t}\alpha_{j}$. Then, the DDIM sampling is done according to $p_{\theta}({\mathbf{z}}_{t-1}|{\mathbf{z}}_{t})=\mathcal{N}({\mathbf{z}}_{t-1}|\bm{\mathbf{\mu}}_{\theta}({\mathbf{z}}_{t},t),\mathbf{\Sigma}(t))$, where $\theta$ are the parameters of the neural network and,
$\mathbf{\Sigma}(t)$ $=\sigma_{t}^{2}{\mathbf{I}}$ (1)
$\bm{\mathbf{\mu}}_{\theta}({\mathbf{z}}_{t},t)$ $=\sqrt{\bar{\alpha}_{t-1}}\left(\frac{{\mathbf{z}}_{t}-\sqrt{1-\bar{\alpha}_{t}}\cdot\bm{\mathbf{\epsilon}}_{\theta}({\mathbf{z}}_{t},t)}{\sqrt{\bar{\alpha}_{t}}}\right)+$
$~{}~{}~{}~{}\sqrt{1-\bar{\alpha}_{t-1}-\sigma^{2}_{t}}\cdot\bm{\mathbf{\epsilon}}_{\theta}({\mathbf{z}}_{t},t).$
As in (Dou & Song, 2024) we fix $\sigma_{t}=\eta\cdot\sqrt{\beta_{t}\cdot\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}}$ with $\eta$ being a hyper-parameter. We denote the approximate posterior mean of $\mathbb{E}[{\mathbf{z}}_{0}|{\mathbf{z}}_{t}]$ by $\bar{{\mathbf{z}}}_{0}({\mathbf{z}}_{t})\coloneqq\frac{1}{\sqrt{\bar{\alpha}_{t}}}({\mathbf{z}}_{t}-\sqrt{1-\bar{\alpha}_{t}}\cdot\bm{\mathbf{\epsilon}}_{\theta}({\mathbf{z}}_{t},t))$ (Robbins, 1956; Efron, 2011; Chung et al., 2023b).
Sequential Monte Carlo (SMC). SMC is an important technique for sampling in probabilistic graphical models in which exact posterior inference is intractable. The SMC breaks the sampling process down to intermediate steps, allowing efficient sampling through a recursive procedure (Doucet et al., 2001; Del Moral et al., 2012; Naesseth et al., 2019; Chopin et al., 2020).
One family of probabilistic models for which SMC is especially known is state-space models (SSMs), also known as Hidden Markov Models (HMMs). In general, the following quantities need to be defined in SSMs, (1) a prior distribution over the initial state $p({\mathbf{z}}_{T})$, (2) a transition distribution that defines the dynamics between states $p({\mathbf{z}}_{t}|{\mathbf{z}}_{t+1})~{}\forall t<T$, and (3) a measurement model $p({\mathbf{y}}_{t}|{\mathbf{z}}_{t})~{}\forall t<T$. The goal is to sample from the posterior distribution $p({\mathbf{z}}_{t:T}|{\mathbf{y}}_{t:T-1})$. To do so, SMC starts by sampling $N$ particles $\{{\mathbf{z}}^{(i)}_{T}\}_{i=1}^{N}$ from the prior distribution. Then, at each step, given the previous particle set $\{{\mathbf{z}}^{(i)}_{t}\}_{i=1}^{N}$ new samples are taken from a proposal distribution ${\mathbf{z}}^{(i)}_{t-1}\sim\pi({\mathbf{z}}^{(i)}_{t-1}|{\mathbf{z}}^{(i)}_{t})~{}\forall i\in\{1,...,N\}$. The particles are then weighted and resampled according to the new proposed sequences $\{{\mathbf{z}}^{(i)}_{t-1:T}\}_{i=1}^{N}$. The proposal distribution serves as an approximation to the posterior distribution. Its support needs to contain the support of the posterior density. The weighting function corrects the approximation by assigning a weight to each particle to adjust its probability. And the resampling step aims to remove unlikely particles according to the model (Särkkä, 2013).
3 Related Work
Inverse problems have a long and evolving history, with methodologies that have undergone significant advances over the years (Daras et al., 2024). Recently, diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021c) have emerged as effective priors for solving inverse problems in image data (Wang et al., 2023; Kawar et al., 2022; Chung et al., 2023b; Dou & Song, 2024; Rout et al., 2023; Song et al., 2023; Sun et al., 2024; Choi et al., 2021; Chung et al., 2023a).
In (Song et al., 2021c) it was shown that to sample from the posterior distribution, $p({\mathbf{x}}_{0}|{\mathbf{y}})$, one can solve a stochastic differential equation based on the prior score, $\nabla_{{\mathbf{x}}_{t}}\log~{}p_{\theta}({\mathbf{x}}_{t})$, and the conditional score, $\nabla_{{\mathbf{x}}_{t}}\log~{}p_{\theta}({\mathbf{y}}|{\mathbf{x}}_{t})$. Although the first term is easy to compute, the latter term requires integration over the full diffusion path from time $t$ to 00. A useful and easy-to-calculate approximation found in several studies is $p_{\theta}({\mathbf{y}}|{\mathbf{x}}_{t})\approx p_{\theta}({\mathbf{y}}|\mathbb{E}[{\mathbf{x}}_{0}|{\mathbf{x}}_{t}])$, which is readily available at each step (Chung et al., 2023b; Song et al., 2023; Wu et al., 2024). Specifically, Diffusion Posterior Sampling (DPS) (Chung et al., 2023b) uses this approximation for linear and non-linear inverse problems with Gaussian and Poisson likelihood models. $\Pi$GDM (Song et al., 2023) introduces pseudoinverse guidance by matching the denoising output and the corrupted image ${\mathbf{y}}$, via transformation of both through a ’pseudoinverse’ of the corruption model. DDNM (Wang et al., 2023) suggested to refine only the contents of the null space during the backward diffusion process. As such it is suited only for linear inverse problems. An additional category of inverse problem approaches that use diffusion models is designed with the objective of asymptotic exactness (Cardoso et al., 2023; Trippe et al., 2023; Wu et al., 2024; Dou & Song, 2024). These methods utilize SMC techniques targeting exact sampling from the posterior distribution $p({\mathbf{x}}_{0}|{\mathbf{y}})$. Specifically, SMC-Diff (Trippe et al., 2023) applies particle filtering for inpainting in motif scaffolding, and MCGDiff (Cardoso et al., 2023) is designed for linear inverse problems only. Hence, both approaches are not suited for inverse problems with latent-space diffusion models. TDS (Wu et al., 2024), a recent SMC-based method, solves general inverse problem tasks using the twisting technique. This method also uses the approximation of DPS, but by applying SMC sampling it can correct for it.
FPS (Dou & Song, 2024) is also a recent method based on SMC with auxiliary variables. FPS generates a sequence of observations ${\mathbf{y}}_{1:T}$ based on a duplex diffusion process, one process at the ${\mathbf{x}}$ space and the other process at the ${\mathbf{y}}$ space. Since this method is designed for linear inverse problems only, it permits tractable Bayesian inference. Our method combines the ideas of both TDS and FPS to obtain the best of both. Namely, we use the posterior mean approximation and ${\mathbf{y}}_{1:T}$ in our SMC sampling process. As we will show, this combination can be helpful in both understanding the general semantics of an image and capturing fine details.
Several inverse sampling methods were specifically tailored for latent diffusion models. PSLD (Rout et al., 2023) extend DPS (Chung et al., 2023b) by incorporating an additional gradient update step to guide the diffusion process to sample latent representations that maintain the integrity of the decoding-encoding transformation, ensuring it remains non-lossy. STSL (Rout et al., 2024) presents a novel sampler with a tractable reverse process using an efficient second-order approximation. Comparative analysis with STSL was not feasible due to the absence of publicly available code, making replication challenging. Resample (Song et al., 2024), a contemporary method alongside PSLD, introduces a strategy for addressing general inverse problems using pretrained latent diffusion models, tackling the complexities posed by encoder and decoder nonlinearity. Resample algorithm includes hard data consistency to obtain latent variable that is consistent with the observed measurements, and then employs a resampling scheme to map the sample back onto the correct noisy data manifold and continue the reverse sampling process. Concurrent to this study Nazemi et al. (2024) proposed a particle filtering approach. Their method builds on PSLD and DPS update in the proposal distribution. Similarly to TDS (Wu et al., 2024) the connection to the labels is only through ${\mathbf{z}}_{0}$ using the approximate mean estimator. Since these methods share commonalities, we compare only to the latter in the experimental section.
[…]
Figure 3: The graphical model of LD-SMC. In gray observed variables and in white are latent variables.
4 Method
Given a corrupted image ${\mathbf{y}}_{0}$, the goal is to sample $\hat{{\mathbf{z}}}_{0}\sim p_{\theta}({\mathbf{z}}_{0}|{\mathbf{y}}_{0})$ using a pre-trained latent diffusion model as prior. Then, we can transform this sample into an image by applying a pre-trained decoder ${\mathcal{D}}$, i.e. ${\mathbf{x}}_{0}\coloneqq{\mathcal{D}}(\hat{{\mathbf{z}}}_{0})$. We first define a generative model for the data and then apply Bayesian inference on all latent variables using blocked Gibbs sampling and SMC. Specifically, we use the forward diffusion process to augment the model with additional auxiliary observations, and then apply posterior inference using a sequential Monte Carlo (SMC) approach based on the backward diffusion process. The corresponding graphical model can be seen in Figure 3.
4.1 The Generative Model
We now explicitly define the data generation model based on the forward diffusion process of DDIM (Song et al., 2021a),
$1.~{}{\mathbf{z}}_{0}\sim p({\mathbf{z}}_{0})$
$2.~{}{\mathbf{z}}_{T}|{\mathbf{z}}_{0}\sim\mathcal{N}(\sqrt{\bar{\alpha}_{T}}{\mathbf{z}}_{0},(1-\bar{\alpha}_{T}){\mathbf{I}})),$
$3.~{}{\mathbf{z}}_{t-1}|{\mathbf{z}}_{t},{\mathbf{z}}_{0}\sim p({\mathbf{z}}_{t-1}|{\mathbf{z}}_{t},{\mathbf{z}}_{0})\quad\forall t\in\{2,...,T\},$
$4.~{}{\mathbf{y}}_{t}|{\mathbf{z}}_{t}\sim\mathcal{N}({\mathcal{A}}(\underbrace{{\mathcal{D}}({\mathbf{z}}_{t}))}_{{\mathbf{x}}_{t}},\tau^{2}{\mathbf{I}})\quad\forall t\in\{0,...,{T-1}\}.$
Here $p({\mathbf{z}}_{0})$ is a prior distribution over ${\mathbf{z}}_{0}$, $p({\mathbf{y}}_{t}|{\mathbf{z}}_{t})$ is defined by the corruption model, and
$p({\mathbf{z}}_{t-1}|{\mathbf{z}}_{t},{\mathbf{z}}_{0})=$ (2)
$\mathcal{N}\left({\mathbf{z}}_{t-1}|\sqrt{\bar{\alpha}_{t-1}}{\mathbf{z}}_{0}+\sqrt{1-\bar{\alpha}_{t-1}-\sigma^{2}_{t}}\cdot\frac{{\mathbf{z}}_{t}-\sqrt{\bar{\alpha}_{t}}{\mathbf{z}}_{0}}{\sqrt{1-\bar{\alpha}_{t}}},\sigma_{t}^{2}{\mathbf{I}}\right),$
is defined by the forward diffusion process. Introducing unobserved data is a known technique in statistics for conducting effective Markov chain Monte Carlo (MCMC) sampling (Van Dyk & Meng, 2001; Dou & Song, 2024). In our case, we can use it while leveraging the dependencies between the variables in order to build an efficient SMC sampling procedure, as described in the next section.
4.2 Sampling Procedure
Given the generative model defined in Section 4.1, our aim is to apply Bayesian inference over the latent variables. In broad strokes, to obtain a sample $\hat{{\mathbf{z}}}_{0}\sim p({\mathbf{z}}_{0}|{\mathbf{y}}_{0})$ we use blocked Gibbs sampling to sample in turns from $p({\mathbf{y}}_{1:T-1}|{\mathbf{z}}_{0:T},{\mathbf{y}}_{0})$ and then use SMC to sample from $p({\mathbf{z}}_{0:T}|{\mathbf{y}}_{0:T-1})$. Specifically, we propose the following procedure:
  1. 1.
    Obtain an initial guess for $\hat{{\mathbf{z}}}_{0}$ (detailed in Sec. 4.2.1),
  2. 2.
    Repeat for some fixed number of steps:
    1. (a)
      Sample, ${\mathbf{z}}_{1:T}\sim p({\mathbf{z}}_{1:T}|\hat{{\mathbf{z}}}_{0},{\mathbf{y}}_{0})=p({\mathbf{z}}_{1:T}|\hat{{\mathbf{z}}}_{0})$ according to the forward process of DDIM (Eq. 2).
    2. (b)
      Sample, ${\mathbf{y}}_{1:T-1}\sim p({\mathbf{y}}_{1:T-1}|\hat{{\mathbf{z}}}_{0},{\mathbf{z}}_{1:T},{\mathbf{y}}_{0})=p({\mathbf{y}}_{1:T-1}|{\mathbf{z}}_{1:T-1})=\prod_{t=1}^{T-1}\mathcal{N}({\mathbf{y}}_{t}|{\mathcal{A}}({\mathcal{D}}({\mathbf{z}}_{t})),\tau^{2}{\mathbf{I}})$,
    3. (c)
      Sample N particles $\{\hat{{\mathbf{z}}}_{T}^{(i)}\}_{i=1}^{N}\sim\mathcal{N}({\mathbf{z}}_{T},\xi^{2}{\mathbf{I}})$,
    4. (d)
      Sample $\hat{{\mathbf{z}}}_{0:T-1}\sim p_{\theta}({\mathbf{z}}_{0:T-1}|{\mathbf{y}}_{0:T-1},\{\hat{{\mathbf{z}}}_{T}^{(i)}\}_{i=1}^{N})$ using SMC based on a pre-trained diffusion model and select one particle $\hat{{\mathbf{z}}}_{0}$ for the next iteration (detailed in Sec. 4.2.2).
Here we use the dynamics of the forward process and the graphical model dependencies in steps (a) and (b). In step (c) we sample from a Gaussian centered at the final ${\mathbf{z}}_{T}$ having a small std $\xi$ instead of the prior distribution as commonly done and described in Section 2. Although empirically we found that this step did not have a significant effect on the final model performance, we did it to be better aligned with the generated ${\mathbf{y}}_{t}$ from the forward process. In addition, we found that one forward-backward pass suffices to achieve good results. Hence, in our empirical evaluations, we apply Step 2 only once. In conclusion, steps (a), (b) and (c) are fully defined. The two steps that are not straightforward are how to obtain an initial guess for $\hat{{\mathbf{z}}}_{0}$ (step 1) and how to perform the sampling process in step (d). We discuss both next.
4.2.1 Initial Guess for $\hat{{\mathbf{z}}}_{0}$
The first challenge is to obtain an initial $\hat{{\mathbf{z}}}_{0}$ (step 1. in the sampling procedure). To reduce the variance in the process and accelerate convergence, we performed the following optimization procedure in pixel space:
$\hat{{\mathbf{x}}}_{0}$ $=\operatorname*{arg\,max}_{\mathbf{x}}log~{}p_{\theta}({\mathbf{y}}_{0}|{\mathbf{x}})=\operatorname*{arg\,min}_{\mathbf{x}}||{\mathbf{y}}_{0}-{\mathcal{A}}({\mathbf{x}})||^{2}_{2},$ (3)
and then we applied the encoder on the outcome, namely $\hat{{\mathbf{z}}}_{0}={\mathcal{E}}(\hat{{\mathbf{x}}}_{0})$. For linear inverse problems, this optimization problem can be solved in closed form (Song et al., 2024; Wang et al., 2023), although it can be costly, as it requires inverting the linear operator ${\mathcal{A}}$. An alternative for this procedure is to apply the optimization process directly in the latent space. However, in our experiments we found that the former option worked better and it did not involve expensive gradient propagation through the decoder.
4.2.2 Posterior Sampling
We now move on to explain step (d) of the sampling procedure. Due to the non-linearity of the decoder, even for linear inverse problems, finding the exact posterior is intractable. One option to overcome this difficulty is to use SMC sampling. In what follows, we first describe an approximate posterior for $p_{\theta}({\mathbf{z}}_{0:T}|{\mathbf{y}}_{0:T-1})$, then we suggest an iterative procedure based on SMC sampling.
First, notice that because of the structure of the model, the posterior density of the r.v. ${\mathbf{z}}_{t}$ at each step $t$ depends only on ${\mathbf{z}}_{t+1:T}$. Hence, only $p_{\theta}({\mathbf{z}}_{t:T}|{\mathbf{y}}_{0:T-1})$ needs to be computed at each time step $t$. However, even computing an unnormalized quantity of that posterior can be costly. Therefore, we make the following assumption $p_{\theta}({\mathbf{z}}_{t:T}|{\mathbf{y}}_{0:T-1})\approx p_{\theta}({\mathbf{z}}_{t:T}|{\mathbf{y}}_{t:T-1},{\mathbf{y}}_{0})$. We assume this as ${\mathbf{y}}_{0}$ stores all the input information. Now we can arrive at the following recursive formula (to prevent cluttered notations we omit here the subscript of $\theta$ from the probability densities):
$p({\mathbf{z}}_{t:T}|{\mathbf{y}}_{t:T},{\mathbf{y}}_{0})$
$\propto p({\mathbf{y}}_{t}|{\mathbf{z}}_{t:T},{\mathbf{y}}_{t+1:T},{\mathbf{y}}_{0})p({\mathbf{z}}_{t:T}|{\mathbf{y}}_{t+1:T},{\mathbf{y}}_{0})$
$=p({\mathbf{y}}_{t}|{\mathbf{z}}_{t})p({\mathbf{z}}_{t}|{\mathbf{z}}_{t+1:T},{\mathbf{y}}_{t+1:T},{\mathbf{y}}_{0})p({\mathbf{z}}_{t+1:T}|{\mathbf{y}}_{t+1:T},{\mathbf{y}}_{0})$
$=p({\mathbf{y}}_{t}|{\mathbf{z}}_{t})p({\mathbf{z}}_{t}|{\mathbf{z}}_{t+1},{\mathbf{y}}_{0})p({\mathbf{z}}_{t+1:T}|{\mathbf{y}}_{t+1:T},{\mathbf{y}}_{0})$
$=p({\mathbf{y}}_{t}|{\mathbf{z}}_{t})\frac{p({\mathbf{y}}_{0}|{\mathbf{z}}_{t})}{p({\mathbf{y}}_{0}|{\mathbf{z}}_{t+1})}p({\mathbf{z}}_{t}|{\mathbf{z}}_{t+1})p({\mathbf{z}}_{t+1:T}|{\mathbf{y}}_{t+1:T},{\mathbf{y}}_{0})$
$\approx\frac{p({\mathbf{y}}_{t}|{\mathbf{z}}_{t})p({\mathbf{y}}_{0}|\bar{{\mathbf{z}}}_{0}({\mathbf{z}}_{t}))}{p({\mathbf{y}}_{0}|\bar{{\mathbf{z}}}_{0}({\mathbf{z}}_{t+1}))}p({\mathbf{z}}_{t}|{\mathbf{z}}_{t+1})p({\mathbf{z}}_{t+1:T}|{\mathbf{y}}_{t+1:T},{\mathbf{y}}_{0}).$
Where, in the first transition we used Bayes rule, in the third transition we used the Markovian assumption, in the forth transition we used Bayes rule again, and in the last transition we make an additional approximation and condition on the posterior mean estimator for both time $t$ and time $t+1$. We define $p({\mathbf{y}}_{0}|\bar{{\mathbf{z}}}_{0}({\mathbf{z}}_{t}))=\mathcal{N}({\mathbf{y}}_{0}|{\mathcal{A}}({\mathcal{D}}(\bar{{\mathbf{z}}}_{0}({\mathbf{z}}_{t}))),(1-\bar{\alpha}_{t}){\mathbf{I}}))$ and similarly for $p({\mathbf{y}}_{0}|\bar{{\mathbf{z}}}_{0}({\mathbf{z}}_{t+1}))$, where as in (Wu et al., 2024) the variance term is taken to be the variance of the forward diffusion process.
Figure 4: Comparison between LD-SMC and baseline methods on Gaussian deblurring of images from the FFHQ dataset.
Now, we can derive an SMC procedure using the proposed posterior. Specifically, we used sequential importance sampling with resampling (SISR). The general prescription (adapted to our setting) is as follows (Särkkä, 2013):
  1. 1.
    Collect N particles as detailed in step (c) in Section 4.2.
  2. 2.
    For $t=T-1,...,0$:
    1. i.
      Draw $\{\tilde{{\mathbf{z}}}_{t}^{(i)}\}_{i=1}^{N}$ samples from a proposal distribution (detailed in Sec. 4.2.3):
      $\tilde{{\mathbf{z}}}_{t}^{(i)}\sim\pi(\tilde{{\mathbf{z}}}_{t}|\hat{{\mathbf{z}}}_{t+1}^{(i)},{\mathbf{y}}_{0:T-1})\quad i=1,...,N.$
    2. ii.
      Compute the proposal weights for all $i\in\{1,...,N\}$ according to:
      $w_{t}^{(i)}$ $\propto\frac{p_{\theta}({\mathbf{y}}_{t}|\tilde{{\mathbf{z}}}_{t}^{(i)})p_{\theta}({\mathbf{y}}_{0}|\bar{{\mathbf{z}}}_{0}(\tilde{{\mathbf{z}}}_{t}^{(i)}))p_{\theta}(\tilde{{\mathbf{z}}}_{t}^{(i)}|\hat{{\mathbf{z}}}_{t+1}^{(i)})}{\pi(\tilde{{\mathbf{z}}}^{(i)}_{t}|\hat{{\mathbf{z}}}_{t+1}^{(i)},{\mathbf{y}}_{0:T-1})p_{\theta}({\mathbf{y}}_{t}|\bar{{\mathbf{z}}}_{0}(\hat{{\mathbf{z}}}_{t+1}^{(i)}))},$
      and then normalize the weights to sum to one.
    3. iii.
      Resample $N$ particles (with replacements) from the discrete distribution: $\hat{{\mathbf{z}}}_{t}^{(i)}\sim\{\tilde{{\mathbf{z}}}_{t}^{(1)},...,\tilde{{\mathbf{z}}}_{t}^{(N)}\}$ with weights $\{w_{t}^{(1)},...,w_{t}^{(N)}\}$.
Here, $p_{\theta}(\tilde{{\mathbf{z}}}_{t}^{(i)}|\hat{{\mathbf{z}}}_{t+1}^{(i)})$ is defined according to Eq. 1 and we assume that the resampling step is performed at each iteration. While the derivation is different, the resulting weighting scheme of LD-SMC bears some resemblance to the weighting of TDS (Wu et al., 2024). The main difference between the two methods is the dependence on the auxiliary variables ${\mathbf{y}}_{1:T}$ in the proposal distribution and resampling weights. Empirically, we observed that this additional conditioning helped to better align the sampling with the corrupted image ${\mathbf{y}}_{0}$ compared to using the posterior mean approximation as in (Chung et al., 2023b) and (Wu et al., 2024). Finally, to select one sample at time $t=0$ for terminating the algorithm in Section 4.2 or the next iteration in the Gibbs sampling procedure, we take the particle that has the highest weight.
Table 1: Quantitative results on $1024$ examples of size $256\times 256$ from ImageNet test set. All methods were evaluated under the same experimental setup using LDM.
Inpainting (Box) Gaussian Deblur Super Resolution (8×8$\times$8 ×) FID ($\downarrow$) NIQE ($\downarrow$) LPIPS ($\downarrow$) FID ($\downarrow$) NIQE ($\downarrow$) LPIPS ($\downarrow$) FID ($\downarrow$) NIQE ($\downarrow$) LPIPS ($\downarrow$) Latent DPS 65.45 7.918 0.407 52.48 6.855 0.383 61.02 6.514 0.439 Latent TDS 65.03 7.872 0.406 50.82 6.695 0.379 58.73 7.157 0.454 Resample 90.32 8.464 0.318 46.45 7.411 0.353 87.65 8.290 0.491 PSLD 79.90 9.268 0.410 79.31 7.972 0.474 78.56 7.000 0.467 LD-SMC (Ours) 51.81 5.103 0.355 52.17 6.789 0.382 59.27 6.423 0.437
4.2.3 Proposal Distribution
The SISR algorithm requires access to $\pi(\tilde{{\mathbf{z}}}_{t-1}|\hat{{\mathbf{z}}}_{t}^{(i)},{\mathbf{y}}_{0:T-1})$, a proposal distribution. The optimal choice in the sense of minimizing the variance of the proposal weights is $\pi(\tilde{{\mathbf{z}}}_{t-1}|\hat{{\mathbf{z}}}_{t:T}^{(i)},{\mathbf{y}}_{0:T-1})=p_{\theta}(\tilde{{\mathbf{z}}}_{t-1}|\hat{{\mathbf{z}}}_{t}^{(i)},{\mathbf{y}}_{0:t})$ (Doucet et al., 2000; Särkkä, 2013). However, it cannot be obtained in closed form. Hence, we design an alternative proposal distribution, which we will now describe.
For clarity, we drop here the index notation of the particles. We define the proposal distribution to be a Gaussian $\pi(\tilde{{\mathbf{z}}}_{t-1}|\hat{{\mathbf{z}}}_{t:T},{\mathbf{y}}_{0:T-1})=\mathcal{N}({\mathbf{m}}_{t},{\mathbf{S}}_{t})$ with parameters:
${\mathbf{S}}_{t}$ $=\sigma_{t}^{2}{\mathbf{I}}$ (4)
${\mathbf{m}}_{t}$ $=\bm{\mathbf{\mu}}_{\theta}(\hat{{\mathbf{z}}}_{t},t)-(\gamma_{t}\nabla_{\hat{{\mathbf{z}}}_{t}}||{\mathbf{y}}_{0}-{\mathcal{A}}({\mathcal{D}}(\bar{{\mathbf{z}}}_{0}(\hat{{\mathbf{z}}}_{t})))||_{2}^{2}$
$\qquad\qquad~{}+\lambda_{t}\nabla_{\bm{\mathbf{\mu}}_{\theta}(\hat{{\mathbf{z}}}_{t},t)}||{\mathbf{y}}_{t}-{\mathcal{A}}({\mathcal{D}}(\bm{\mathbf{\mu}}_{\theta}(\hat{{\mathbf{z}}}_{t},t)))||_{2}^{2}).$
Here we set the variance to be the variance of the prior diffusion model; however, other choices are applicable as well. The idea behind our proposal mean is to correct the prior mean estimation by shifting it towards latents that agree more strongly with both ${\mathbf{y}}_{t}$ and ${\mathbf{y}}_{0}$. The second term can be seen as making one gradient update step starting from the current prior mean location.
The $\gamma_{t}$ and $\lambda_{t}$ parameters control the effect of the correction terms to the prior mean. In practice, it is challenging to control the trade-off between the two correction terms. Hence, during the sampling process we start only with the first term and then at some predefined point, $t=s$, we switch to the second term. The intuition here is that during the initial sampling steps the quality of the labels ${\mathbf{y}}_{t}$ may not be good. Therefore, we rely on the first term through the posterior mean estimator to capture the general semantics of the image. However, in later sampling stages the quality of the labels increases (see Figure 5 in the Appendix) and the latter correction term can add fine details to the image. This intuition also relates to the three-stage phenomenon in the diffusion sampling process witnessed in the literature (Yu et al., 2023). Setting $\lambda_{t}=0$ for all time steps reduces the LD-SMC proposal update to that of TDS (Wu et al., 2024). We show in Appendix B an instantiation of $\gamma_{t}$ and $\lambda_{t}$ used in this study.
5 Experiments
Table 2: Quantitative results on $1024$ examples of size $256\times 256$ from FFHQ test set. All methods were evaluated under the same experimental setup using LDM.
Inpainting (Box) Gaussian Deblur Super Resolution (8×8$\times$8 ×) FID ($\downarrow$) NIQE ($\downarrow$) LPIPS ($\downarrow$) FID ($\downarrow$) NIQE ($\downarrow$) LPIPS ($\downarrow$) FID ($\downarrow$) NIQE ($\downarrow$) LPIPS ($\downarrow$) Latent DPS 39.81 7.592 0.236 31.81 6.813 0.285 29.64 6.412 0.282 Latent TDS 39.57 7.602 0.236 33.19 6.879 0.288 30.45 6.411 0.284 Resample 86.79 7.142 0.230 39.80 7.441 0.275 59.23 7.307 0.356 PSLD 47.51 7.480 0.312 36.31 6.802 0.341 40.33 6.803 0.347 LD-SMC (Ours) 37.14 7.520 0.224 32.18 6.566 0.280 30.37 6.456 0.284
5.1 Experimental Setting
We evaluated LD-SMC on ImageNet (Russakovsky et al., 2015) and FFHQ (Karras et al., 2019); both are common in the literature of inverse problems (e.g., (Chung et al., 2023b; Dou & Song, 2024)). In ImageNet samples were conditioned on the class label. The guidance scale was fixed to $1.0$ in all our experiments. Results can be improved by adjusting it (Rombach et al., 2022). Images were resized to $3\times 256\times 256$ and normalized to the range $[0,1]$. We used the latent diffusion model (LDM) VQ-4 (Rombach et al., 2022) for the prior model with the DDIM diffusion sampler (Song et al., 2021a), according to the data split in (Esser et al., 2021). We sampled $1024$ random images from the validation set of each dataset which were used to evaluate all methods. We followed the protocol of (Song et al., 2024) and added Gaussian noise with zero mean and standard deviation $\tau=0.01$ to the corrupted images. Full experimental details are provided in Appendix A.
Compared methods. We compared LD-SMC with several recent SoTA baseline methods; the first two methods were designed for general inverse problems, but were evaluated using pixel space diffusion models, and the latter two were designed specifically for inverse problems in the latent space. Nevertheless, all methods were evaluated under a similar experimental setup using latent diffusion to ensure fairness in the comparisons. (1) Diffusion Posterior Sampling (DPS) (Chung et al., 2023b), which introduces correction to the sampling process of the diffusion through the posterior mean estimator; (2) Twisted Diffusion Sampling (TDS) (Wu et al., 2024), which uses the twisting technique for approximate sequential Monte Carlo sampling; (3) Resample (Song et al., 2024), which applies an optimization procedure during the sampling process to match the approximate posterior mean to the label, and then performs resampling; (4) Posterior Sampling with Latent Diffusion (PSLD) (Rout et al., 2023), which introduces a correction term to the DPS step to “glue” ${\mathbf{z}}_{0}$. For our method and TDS we used $N=5$ particles.
Evaluation metrics. We report the following metrics in the main text, FID (Heusel et al., 2017), NIQE (Mittal et al., 2012), and LPIPS (Zhang et al., 2018). Full results with the PSNR and SSIM metrics (Wang et al., 2004) are deferred to Appendix D. The first two are considered perceptual metrics, lower values in them indicate higher perceptual quality. The other metrics are considered as distortion metrics, which quantify some discrepancy between the generated images and the ground-truth values. Since perceptual metrics and distortion metrics can be in conflict with each other (Blau & Michaeli, 2018), we put more emphasis on perceptual quality. Hence, for all methods, we performed grid search over hyper-parameters and chose the best configuration according to the FID.
5.2 Experimental Results
Quantitative results are shown in Tables 1 and 2. From the tables, LD-SMC is usually the best or second best among all the comparisons. Specifically, on inpainting where extrapolation is needed inside the box and details should be preserved outside the box, LD-SMC can greatly improve over baseline methods, improving the FID score by up to $\sim 13$ points. This property is also manifested in Figure 2 and Figure 6 in the appendix. LD-SMC manages to produce plausible reconstructions while maintaining fine details. This is in contrast to baseline methods, which mainly rely on the posterior mean approximation for reconstruction. The differences are especially highlighted on the ImageNet dataset which has more diversity in it. Also, as is clear from the figures, Resample images suffer from significant artifacts. We speculate that it partly stems from the complete optimization process performed in every few sampling steps according to this method. We observe that this method is well suited for some tasks, such as Gaussian deblurring, but does not perform well on others such as box inpainting.
In Figure 4 we show qualitative results for Gaussian deblurring of images by LD-SMC and baseline methods. From the figure, all methods are able to generate plausible reconstructions on this task. Additional quantitative and qualitative results can be found in Appendix D and E .
5.3 Analysis
Recall that in our proposal update we suggested correcting the prior mean using two terms, one that involves ${\mathbf{y}}_{0}$ and another that involves ${\mathbf{y}}_{t}$. In practice during the sampling process, we first use the former one and then switch to the latter at some fixed time step $s$. This is a hyperparameter of our approach. Table 3 compares the FID and PSNR on ImageNet inpainting task as a function of the diffusion step $s$ in which the switch is made. From the table, when the switch is done at earlier stages of the sampling process, the FID improves, but at the same time, the PSNR degrades. We chose to use $s=333$ since it balances well between the two metrics while giving more emphasis to the perceptual quality.
Table 3: Tradeoff between the proposal update correction terms for $T=1000$ steps.
FID ($\downarrow$) PSNR ($\uparrow$)
$s=0$ 65.76 19.61
$s=T/6$ 61.09 19.25
$s=T/3$ 51.81 18.87
$s=T/2$ 48.77 18.59
6 Limitations
Although our approach has strong empirical results, one limitation of our approach is related to computational demand. The sampling time and the memory demand increase with the number of particles. In addition, compared to TDS, in the resampling step, we need to use the decoder one more time to compute $p({\mathbf{y}}_{t}|{\mathbf{z}}_{t})$ (only forward pass), which can also affect the sampling time. This effect can be mitigated by taking fewer particles or by parallelizing LD-SMC between GPUs. Furthermore, for more challenging tasks like box inpainting, competing algorithms tend to exhibit noticeable artifacts, which limits their applicability.
7 Conclusion
In this study, we presented LD-SMC, a novel method for solving inverse problems in the latent space of diffusion models using SMC. Specifically, we leveraged the forward process of the diffusion process to augment the model with auxiliary observations, one per each timestep, and used these observations to guide the sampling process as part of the backward diffusion process. This framework can be seen as applying one step of blocked Gibbs sampling. To perform SMC sampling, we suggested a novel weighing scheme and a novel proposal distribution. Both are based on information from the auxiliary labels and the true label ${\mathbf{y}}_{0}$. Empirically, we validated LD-SMC against strong baseline methods on common benchmarks. The results suggest that LD-SMC can improve the performance over baseline methods, especially in cases where extrapolation is needed (e.g., in inpainting).
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Full Experimental Details
The experiments were mainly carried out using an NVIDIA A100 having 40GB and 80GB memory. In all experiments, we used the DDIM formulation (Song et al., 2021a), although LD-SMC can be applied with other sampling procedures. For all methods, we performed a hyperparameter search on $\eta\in\{0.05,0.5,1.0\}$. For DPS and TDS we examined several scaling coefficients for the prior mean update, including the ones proposed in each corresponding paper, and found that our proposed update works better for both. For all three methods (LD-SMC, DPS, and TDS) we searched for $\nu_{1}\in\{0,1\}$ and $\kappa_{1}\in\{0.4,0.5,1.0,2.0\}$. For our method, we also performed a grid search over the timestep for the switch between the two terms in the proposal distribution $s\in\{0,250,333\}$. For PSLD, in most cases, the default hyper-parameters suggested in the paper and code didn’t yield good results. Hence, we performed a grid search over PSLD’s hyper-parameters $\gamma_{t}\in\{1e-4,1e-3,1e-2,0.1,0.2\}$ and $\eta_{t}\in\{0.05,0.1,0.2,0.9\}$. For Resample, we found that using $\eta=0.0$, the default value in the code, usually performs the best. Also, we performed a grid search over $\gamma$, the scaling coefficient of the resampling step std in $\{4,8,16,40,80,200,400\}$. For each method, we evaluated visually and using the FID on a sample of images and then picked the best hyperparameter configuration. Then, we sampled $1024$ images using the best configuration. Similarly to Resample, we found that applying an optimization process at the end of the sampling process in the latent space can sometimes improve visibility and metric values. We evaluated all models with and without the final optimization process and picked the best one according to the FID. The optimization procedure was not applied to the inpainting task since it created non-smooth changes at the boundaries of the box, making the images look non-natural.
Figure 5: Evolution of ${\mathbf{y}}_{t}$ over time for different tasks according to forward process of DDIM.
Appendix B Proposal Distribution Scaling Coefficients
Recall that our proposal distribution (Eq. 4 in the main text) is made of two elements. These elements are scaled by two coefficients, $\gamma_{t}$ and $\lambda_{t}$. Here, we provide an explicit formula for these coefficients. We found that our proposed scaling works better than common procedures used in the literature. For consistency with baseline methods, we also used our proposed scaling approach for DPS and TDS, since these methods apply a similar update rule. We tried to use it for PSLD and Resample but it didn’t work well for these baselines.
Let ${\mathbf{g}}_{t}^{1}\coloneqq\nabla_{{\mathbf{z}}_{t}}||{\mathbf{y}}_{0}-{\mathcal{A}}({\mathcal{D}}(\bar{{\mathbf{z}}}_{0}({\mathbf{z}}_{t})))||_{2}^{2}$, and ${\mathbf{g}}_{t}^{2}\coloneqq\nabla_{\bm{\mathbf{\mu}}_{\theta}(\hat{{\mathbf{z}}}_{t},t)}||{\mathbf{y}}_{t}-{\mathcal{A}}({\mathcal{D}}(\bm{\mathbf{\mu}}_{\theta}(\hat{{\mathbf{z}}}_{t},t)))||_{2}^{2}$. We set the scaling coefficients $\gamma_{t}$ and $\lambda_{t}$ the same and according to the following scheme: $\gamma_{t}=\kappa_{1}\cdot(1-\bar{\alpha}_{t})^{\nu_{1}}\cdot\frac{1}{max(||{\mathbf{g}}_{t}^{1}||_{2}^{2},1)}$, and similarly $\lambda_{t}=\kappa_{2}\cdot(1-\bar{\alpha}_{t})^{\nu_{2}}\cdot\frac{1}{max(||{\mathbf{g}}_{t}^{2}||_{2}^{2},1)}$. Here $\{\nu_{1},\nu_{2}\}$ are hyperparameters that controls the effect of the variance scaling, and $\{\kappa_{1},\kappa_{2}\}$ scale the entire term. In practice, since we use either ${\mathbf{g}}_{t}^{1}$ or ${\mathbf{g}}_{t}^{2}$, but not both, we have only one set of hyperparameters which is used for both $\gamma_{t}$ and $\lambda_{t}$. That is, $\nu_{1}\coloneqq\nu_{2}$ and $\kappa_{1}\coloneqq\kappa_{2}$.
Appendix C Forward Process
In Figure 5 we present the evolution of the auxiliary labels ${\mathbf{y}}_{t}$ over time as part of the forward process according to our proposed sampling procedure in Section 4.2, steps (a) & (b). From the figure, we observe a gradual cleaning of noise in the auxiliary labels when advancing from time $t=999$ to time $t=0$.
Appendix D Full Results
Table 4: ImageNet. Box in-painting on $1024$ test examples.
Perceptual Quality Distortion
FID ($\downarrow$) NIQE ($\downarrow$) PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$)
Latent DPS 65.45 7.918 19.19 0.623 0.407
Latent TDS 65.03 7.872 19.21 0.623 0.406
Resample 90.32 8.464 18.16 0.695 0.318
PSLD 79.90 9.268 17.48 0.583 0.410
LD-SMC (Ours) 51.81 5.103 18.87 0.599 0.355
Table 5: ImageNet. Gaussian debluring on $1024$ test examples.
Perceptual Quality Distortion
FID ($\downarrow$) NIQE ($\downarrow$) PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$)
Latent DPS 52.48 6.855 23.61 0.615 0.383
Latent TDS 50.82 6.695 23.57 0.614 0.379
Resample 46.45 7.411 24.36 0.639 0.353
PSLD 79.31 7.972 21.38 0.483 0.474
LD-SMC (Ours) 52.17 6.789 23.60 0.614 0.382
Table 6: ImageNet. Super Resolution (8×8$\times$8 ×) on $1024$ test examples.
Perceptual Quality Distortion
FID ($\downarrow$) NIQE ($\downarrow$) PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$)
Latent DPS 61.02 6.514 21.65 0.523 0.439
Latent TDS 58.73 7.157 21.45 0.515 0.454
Resample 87.65 8.290 22.05 0.532 0.491
PSLD 78.56 7.000 21.54 0.516 0.467
LD-SMC (Ours) 59.27 6.423 21.64 0.521 0.437
Table 7: FFHQ. Box in-painting on $1024$ test examples.
Perceptual Quality Distortion
FID ($\downarrow$) NIQE ($\downarrow$) PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$)
Latent DPS 39.81 7.592 24.15 0.814 0.236
Latent TDS 39.57 7.602 24.24 0.814 0.236
Resample 86.79 7.142 19.75 0.815 0.230
PSLD 47.51 7.480 22.70 0.722 0.312
LD-SMC (Ours) 37.14 7.520 24.08 0.817 0.224
Table 8: FFHQ. Gaussian debluring on $1024$ test examples.
Perceptual Quality Distortion
FID ($\downarrow$) NIQE ($\downarrow$) PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$)
Latent DPS 31.81 6.813 26.25 0.709 0.285
Latent TDS 33.19 6.879 26.13 0.705 0.288
Resample 39.80 7.441 28.45 0.763 0.275
PSLD 36.31 6.802 24.02 0.633 0.341
LD-SMC (Ours) 32.18 6.566 26.60 0.721 0.280
Table 9: FFHQ. Super resolution (8×8$\times$8 ×) on $1024$ test examples.
Perceptual Quality Distortion
FID ($\downarrow$) NIQE ($\downarrow$) PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$)
Latent DPS 29.64 6.412 25.48 0.701 0.282
Latent TDS 30.45 6.412 25.38 0.698 0.284
Resample 59.23 7.307 25.55 0.661 0.356
PSLD 40.33 6.803 23.66 0.615 0.347
LD-SMC (Ours) 30.37 6.456 25.42 0.698 0.284
Appendix E Image Reconstructions
Figure 6: Comparison between LD-SMC and baseline methods on inpainting of images from the FFHQ dataset.
Figure 7: Gaussian debluring. LD-SMC reconstruction of images from FFHQ (left) and ImageNet (right).
Figure 8: Super resolution. LD-SMC reconstruction of images from FFHQ (left) and ImageNet (right).